OpenStudy (anonymous):

medal !! Find the length of the radius of the following circle. (x + 2)2 + y2 = 10 5 √(10) 100

OpenStudy (therealmeeeee):

The rule for the equation of a circle is x^2+y^2=r^2. So if you have x^2+y^2=36, just squareroot the 36 and you'll have your radius. In this case, the radius is 6.

OpenStudy (therealmeeeee):

did this help u

OpenStudy (anonymous):

so its 5

OpenStudy (therealmeeeee):

yes

OpenStudy (therealmeeeee):

no 6

OpenStudy (therealmeeeee):

yeah 5 i mean

OpenStudy (anonymous):

okay i get it thanks :) medal for you:)

OpenStudy (therealmeeeee):

ty

OpenStudy (anonymous):

could you help me with this one Find the coordinates of the center of the following circle. (x + 3)2 + (y - 6)2 = 24 (-3,6) (3,-6) (6,-3)

OpenStudy (therealmeeeee):

ok Equation of a circle with center at (h, k) id given by (x - h)² + (y - k)² = r², where r is the radius of the circle Now, (x + 3)² + (y - 6)² = 24 shows that the center is at (h, k) = (- 3, 6) .

OpenStudy (anonymous):

Thanks :D

OpenStudy (therealmeeeee):

nps